The sum of Mila's and Johny's ages is 36 years. In 5 difference of their ages will be 12 years. What are their ages now?
Answer by chndnbl(31)

Let Mila's and Johny' be x and y.
From the first condition:
x + y = 36
from the second condition:
x - y = 12
by using elimination method,
2x = 24
x = 12 years and
y = 36 - 12=24 years